the intensity of illumination from a source of light varies inversely as the square of the distance from the source. (a) at what distance from a source of light will the intensity of illumination be half as great as it is at a distance of 6 inches? (b) how much closer to a source of light must an object be moved so that the intensity of illumination which it receives will be 44 per cent more than that which it receives at a distance of 3 feet?
Answer by rothauserc(4718)

(a)intensity of illumination from a source of light at 6 inches is 1/6^2 = 1/36
:
half the illumination is (1/36)/2 = 1/72
:
square root of 72 = 8.49 inches
:
(b) at three feet, illumination is 1/9
:
(1/9) + 0.44 * (1/9) = 0.16
:
square root(0.16) = 0.4 feet
:
3 feet - 0.4 feet = 2.6 feet closer
: